# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ Appendix E $-\ell$ Converse Analysis for Near-Constant Weight Designs
$M_{j}^{\prime}|(\mathbf{T}_{s},\mathbf{T}_{1},...,\mathbf{T}_{j-1})\sim\mathrm{Hg}((1+o(1))k\Delta,(1+o(1))e^{-\nu}k\Delta,\Delta).$ (185)
The Chernoff bound and matching anti-concentration for the hypergeometric distribution (see Lemma 6 in Appendix G) gives that
$\mathbb{P}\Big{(}M_{j}^{\prime}=\frac{Cn\nu e^{-\nu}}{k}\big{|}(\mathbf{T}_{s},\mathbf{T}_{1},...,\mathbf{T}_{j-1})\Big{)}=\exp\Big{(}-\frac{\nu n}{k}\big{[}D(Ce^{-\nu}\|e^{-\nu})+o(1)\big{]}\Big{)}$ (186)
@@
Bounding $\mathbb{P}(j\in\mathcal{K}^{\prime}_{C,\zeta})$ from below: Under $M_{j}^{\prime}=\frac{Cn\nu e^{-\nu}}{k}$, with the randomness of $\mathbf{Z}$ we have $M_{j0}^{\prime}\sim\text{Bin}\big{(}\frac{Cne^{-\nu}\nu}{k},\rho\big{)}$, which should equal $\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}$ to ensure $j\in\mathcal{K}^{\prime}_{C,\zeta}$. Specifically, for $1\leq j\leq k^{\xi}$ we proceed as follows:
$\mathbb{P}\big{(}j\in\mathcal{K}^{\prime}_{C,\zeta}\big{|}(\mathbf{T}_{s},\mathbf{T}_{1},...,\mathbf{T}_{j-1})\big{)}$ (187)
$\quad=\mathbb{P}\Big{(}M_{j}^{\prime}=\frac{Ce^{-\nu}\nu n}{k},M_{j0}^{\prime}=\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}\big{|}(\mathbf{T}_{s},\mathbf{T}_{1},...,\mathbf{T}_{j-1})\Big{)}$ (188)
$\quad=\mathbb{P}\Big{(}M_{j}^{\prime}=\frac{Cn\nu e^{-\nu}}{k}\big{|}(\mathbf{T}_{s},\mathbf{T}_{1},...,\mathbf{T}_{j-1})\Big{)}\cdot\mathbb{P}\Big{(}\text{Bin}\Big{(}\frac{C\nu e^{-\nu}n}{k},\rho\Big{)}=\frac{\zeta\cdot C\nu e^{-\nu}n}{k}\Big{)}$ (189)
$\quad\geq\exp\left(-\frac{\nu ne^{-\nu}}{k}\Big{[}e^{\nu}D\big{(}Ce^{-\nu}\|e^{-\nu}\big{)}+C\cdot D(\zeta\|\rho)+o(1)\Big{]}\right)$ (190)
$\quad:=\exp\Big{(}-\frac{\nu ne^{-\nu}}{k}\big{[}f_{1}(C,\zeta,\rho,\nu)+o(1)\big{]}\Big{)}:=P_{\text{in}},$ (191)
where in (190) we substitute (186) and apply (26) in Lemma 1 (with leading factor absorbed into the exponent), and in (191) we define
$f_{1}(C,\zeta,\rho,\nu)=e^{\nu}D\big{(}Ce^{-\nu}\|e^{-\nu}\big{)}+C\cdot D(\zeta\|\rho).$ (192)
@@
$k^{\prime}_{C,\zeta}=\sum_{j=1}^{k}\mathbbm{1}(j\in\mathcal{K}^{\prime}_{C,\zeta})\geq\sum_{j=1}^{k^{\xi}}\mathbbm{1}(j\in\mathcal{K}^{\prime}_{C,\zeta}):=k^{\prime}_{C,\zeta,\xi},$ (193)
and from (191) we have $(k^{\prime}_{C,\zeta,\xi}|\mathbf{T}_{s})$ stochastically dominates $\text{Bin}(k^{\xi},P_{\text{in}})$, and thus
$\mathbb{P}\big{(}k^{\prime}_{C,\zeta,\xi}\leq z\,\big{|}\,\mathbf{T}_{s}\big{)}\leq\mathbb{P}\big{(}\text{Bin}(k^{\xi},P_{\text{in}})\leq z\big{)}$ (194)
for any $z$. Combining this property with (183) and (193), we can proceed as
$\mathbb{P}\big{(}k_{C,\zeta}=0\,|\,\mathbf{T}_{s}\big{)}$ $\leq\mathbb{P}\Big{(}k_{C,\zeta}^{\prime}\leq C_{0}\log k\big{|}\mathbf{T}_{s}\Big{)}$ (195)
$\leq\mathbb{P}\Big{(}k_{C,\zeta,\xi}^{\prime}\leq C_{0}\log k\big{|}\mathbf{T}_{s}\Big{)}$ (196)
$=\mathbb{P}\Big{(}\text{Bin}(k^{\xi},P_{\text{in}})\leq C_{0}\log k\Big{)},$ (197)
@@
$\mathbb{P}\big{(}k_{C,\zeta}=0\big{)}\leq\mathbb{P}\Big{(}\text{Bin}(k^{\xi},P_{\text{in}})\leq C_{0}\log k\Big{)}.$ (198)
We now claim that to ensure that (C1) holds with $1-o(1)$ probability, it suffices to ensure
$k^{\xi}P_{\text{in}}=\exp\Big{(}\xi\log k-\frac{\nu ne^{-\nu}[f_{1}(C,\zeta,\rho,\nu)+o(1)]}{k}\Big{)}\geq\exp\Big{(}\eta_{1}\log\frac{p}{k}\Big{)}\to\infty$ (199)
for some $\eta_{1}>0$. Indeed, if (199) holds, then by the Chernoff bound (see (24) in Lemma 1), we have
$\mathbb{P}\Big{(}\text{Bin}(k^{\xi},P_{\text{in}})\leq C_{0}\log k\Big{)}$ $\leq\exp\Big{(}-k^{\xi}P_{\text{in}}\Big{[}\frac{C_{0}\log k}{k^{\xi}P_{\text{in}}}\log\Big{(}\frac{C_{0}\log k}{k^{\xi}P_{\text{in}}}\Big{)}+\frac{C_{0}\log k}{k^{\xi}P_{\text{in}}}-1\Big{]}\Big{)}$ (200)
$\leq\exp\Big{(}-\frac{1}{2}k^{\xi}P_{\text{in}}\Big{)},$ (201)
where the second line holds since $\frac{C_{0}\log k}{k^{\xi}P_{\text{in}}}\leq\frac{C_{0}\log k}{\exp(\eta_{1}\log\frac{p}{k})}=o(1)$ for large enough $k$. The claim thus follows by combining with (198).
It remains to deduce the threshold from (199). By $k=\Theta(p^{\theta})$ we have $\log k\sim\frac{\theta}{1-\theta}\log\frac{p}{k}$, and then by choosing $\xi$ arbitrarily close to $1$ we can ensure (199) by
$n\leq(1-\eta_{1})\frac{\frac{\theta}{1-\theta}k\log\frac{p}{k}}{\nu e^{-\nu}f_{1}(C,\zeta,\rho,\nu)}$ (202)
for some $\eta_{1}>0$ possibly different from the one in (199).
@@
The subsequent analysis is built on condition (C1) that ensures $k_{C,\zeta}\geq 1$, and we consider a single specific index $j\in\mathcal{K}_{C,\zeta}$. We will study the placement of each non-defective item $j^{\prime}$ for $k+1\leq j^{\prime}\leq p$. For the given $j\in\mathcal{K}_{C,\zeta}$ and a fixed $k+1\leq j^{\prime}\leq p$, we define $R_{j^{\prime}}=|\mathcal{N}_{0,j^{\prime}}\cup\mathcal{M}_{j,j^{\prime}}|$ as the number of tests in $\mathcal{N}_{0}\cup\mathcal{M}_{j}$ that contain item $j^{\prime}$, and define $R_{j^{\prime}}^{\prime}$ as the number of placements of item $j^{\prime}$ that are in $\mathcal{N}_{0}\cup\mathcal{M}_{j}$. It is evident that $R_{j^{\prime}}^{\prime}\geq R_{j^{\prime}}$ always holds, and $R_{j^{\prime}}^{\prime}>R_{j^{\prime}}$ when item $j^{\prime}$ is placed more than once into some test in $\mathcal{N}_{0}\cup\mathcal{M}_{j}$. To formulate condition (C2), we are primarily interested in $R_{j^{\prime}}$, but it is more convenient to first study $R_{j^{\prime}}^{\prime}$. Note that under the event $\mathscr{A}_{3}$ in (180) which states
$ne^{-\nu}(1+o(1))=N_{0}\leq|\mathcal{N}_{0}\cup\mathcal{M}_{j}|\leq N_{0}+\ell\Delta=ne^{-\nu}(1+o(1)),$ (203)
a placement of item $j^{\prime}$ increments $R_{j^{\prime}}^{\prime}$ by $1$ with probability $e^{-\nu}(1+o(1))$, thus we have $R_{j^{\prime}}^{\prime}\sim\text{Bin}(\Delta,e^{-\nu}(1+o(1)))$ with respect to the randomness in placing item $j^{\prime}$ into tests.
Bounding the difference between $R_{j^{\prime}}$ and $R_{j^{\prime}}^{\prime}$: We are ultimately interested in $R_{j^{\prime}}$, but we have characterized $R_{j^{\prime}}^{\prime}$ as an intermediate step. Accordingly, we proceed to quantify the difference between $R_{j^{\prime}}$ and $R_{j^{\prime}}^{\prime}$. We envision that the $\Delta$ placements of item $j^{\prime}$ are done subsequently, and let $D_{j^{\prime}}$ be the number of “collisions”, defined as
$D_{j^{\prime}}=\sum_{i=1}^{\Delta}\mathbbm{1}\big{(}\text{the }i\text{-th placement coincides with the }i_{1}\text{-th placement for some }i_{1}<i\big{)}.$ (204)
@@
$R^{\prime}_{j^{\prime}}-D_{j^{\prime}}\leq R_{j^{\prime}}\leq R^{\prime}_{j^{\prime}},$ (205)
@@
$\mathbb{P}\Big{(}D_{j^{\prime}}\geq D_{0}\Big{)}\leq\binom{\Delta}{D_{0}}\Big{(}\frac{\Delta}{n}\Big{)}^{D_{0}}\leq\Big{(}\frac{e\Delta^{2}}{D_{0}n}\Big{)}^{D_{0}}.$ (206)
Thus, we take $D_{0}=\frac{10}{\theta}$ in (206) and apply a union bound over $k+1\leq j^{\prime}\leq p$, yielding
$\mathbb{P}\Big{(}\sup_{k+1\leq j^{\prime}\leq p}D_{j^{\prime}}\geq\frac{10}{\theta}\Big{)}$ $\leq p\cdot\Big{(}\frac{e\theta\Delta^{2}}{10n}\Big{)}^{\frac{10}{\theta}}$ (207)
$=\exp\Big{(}\log p+\frac{10}{\theta}\log\Big{(}\frac{e\theta\nu^{2}n}{10k}\Big{)}-\frac{10}{\theta}\log k\Big{)}$ (208)
$\leq k^{-5},$ (209)
where (209) holds when $k$ is large enough because $p=\Theta(k^{1/\theta})$ and $n=\Theta(k\log k)$. In subsequent analyses, we implicitly suppose (via a union bound) that we are on the event
$D_{j^{\prime}}<\frac{10}{\theta},\leavevmode\nobreak\ \forall\leavevmode\nobreak\ k+1\leq j^{\prime}\leq p,$ (210)
@@
Studying the condition for (176): Given $R_{j^{\prime}}$, the randomness of the noise $\mathbf{Z}$ gives
$G_{j,j^{\prime},1}\sim\text{Bin}(R_{j^{\prime}},\rho),\leavevmode\nobreak\ \leavevmode\nobreak\ G_{j,j^{\prime},2}=R_{j^{\prime}}-G_{j,j^{\prime},1}.$ (211)
Thus, (176) can be expressed as $2G_{j,j^{\prime},1}-R_{j^{\prime}}>(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}$, or equivalently
$G_{j,j^{\prime},1}>\Big{(}\frac{1}{2}-\zeta\Big{)}\frac{Cn\nu e^{-\nu}}{k}+\frac{1}{2}R_{j^{\prime}}.$ (212)
@@
$\mathbb{P}\Big{(}G_{j,j^{\prime},1}>\big{(}\frac{1}{2}-\zeta\big{)}\frac{Cn\nu e^{-\nu}}{k}+\frac{1}{2}R_{j^{\prime}}\Big{)}$ (214)
$\geq\mathbb{P}\Big{(}R_{j^{\prime}}^{\prime}=\frac{dn\nu e^{-\nu}}{k}\Big{)}\mathbb{P}\Big{(}\text{Bin}(R_{j^{\prime}},\rho)>\big{(}\frac{1}{2}-\zeta\big{)}\frac{Cn\nu e^{-\nu}}{k}+\frac{1}{2}R_{j^{\prime}}\Big{|}R_{j^{\prime}}^{\prime}=\frac{dn\nu e^{-\nu}}{k}\Big{)}$ (215)
@@
$\quad\quad\quad\quad\cdot\mathbb{P}\Big{(}\text{Bin}\big{(}\frac{(1+o(1))dn\nu e^{-\nu}}{k},\rho\big{)}>\frac{d+C(1-2\zeta)}{2}\frac{n\nu e^{-\nu}}{k}\Big{)}$ (216)
$\geq\exp\left(-\frac{\nu ne^{-\nu}}{k}\Big{[}e^{\nu}\cdot D(de^{-\nu}\|e^{-\nu})+d\cdot D\Big{(}\frac{1}{2}+\frac{C(1-2\zeta)}{2d}\big{\|}\rho\Big{)}+o(1)\Big{]}\right)$ (217)
$:=\exp\Big{(}-\frac{\nu ne^{-\nu}}{k}\big{[}g(C,\zeta,d,\rho,\nu)+o(1)\big{]}\Big{)},$ (218)
where in (216) we use $R_{j^{\prime}}^{\prime}\sim\text{Bin}(\Delta,e^{-\nu}+o(1))$ along with the event (210) and $R_{j^{\prime}}^{\prime}=\frac{dn\nu e^{-\nu}}{k}$, from which (205) gives $R_{j^{\prime}}=\frac{dn\nu e^{-\nu}(1+o(1))}{k}$ (since $\frac{n}{k}\to\infty$ and $\frac{10}{\theta}$ is a constant); then in (217) we apply (26) in Lemma 1 to the two probability terms (with leading factors absorbed into the exponent), and in (218) we introduce the following shorthand:
$g(C,\zeta,d,\rho,\nu)=e^{\nu}\cdot D\big{(}de^{-\nu}\|e^{-\nu}\big{)}+d\cdot D\Big{(}\frac{1}{2}+\frac{C(1-2\zeta)}{2d}\big{\|}\rho\Big{)}.$ (219)
@@
$g(C,\zeta,d,\rho,\nu)$ $=d\log d+(e^{\nu}-d)\log\frac{1-de^{-\nu}}{1-e^{-\nu}}+\frac{d+A}{2}\log\frac{d+A}{2d\rho}+\frac{d-A}{2}\log\frac{d-A}{2d(1-\rho)}.$ (220)
Optimizing $d$: Our next step is to specify $d$ that minimizes $g(C,\zeta,d,\rho,\nu)$ to render the tightest lower bound (218). Note that range of $d$ is given by $|A|\leq d\leq e^{\nu}$ in (213), and we define
$d^{*}=\text{arg}\min\leavevmode\nobreak\ g(C,\zeta,d,\rho,\nu),\leavevmode\nobreak\ \leavevmode\nobreak\ \text{subject to }|A|\leq d\leq e^{\nu}.$ (221)

### 23 Oct 2024 v2

@@ Appendix E $-\ell$ Converse Analysis for Near-Constant Weight Designs
$M_{j}^{\prime}|(\mathbf{T}_{s},\mathbf{T}_{1},...,\mathbf{T}_{j-1})\sim\mathrm{Hg}((1+o(1))k\Delta,(1+o(1))e^{-\nu}k\Delta,\Delta).$ (E.10)
The Chernoff bound and matching anti-concentration for the hypergeometric distribution (see Lemma G.2 in Appendix G) then gives that
$\mathbb{P}\Big{(}M_{j}^{\prime}=\frac{Cn\nu e^{-\nu}}{k}\big{|}(\mathbf{T}_{s},\mathbf{T}_{1},...,\mathbf{T}_{j-1})\Big{)}=\exp\Big{(}-\frac{\nu n}{k}\big{[}D(Ce^{-\nu}\|e^{-\nu})+o(1)\big{]}\Big{)}$ (E.11)
@@
Bounding $\mathbb{P}(j\in\mathcal{K}^{\prime}_{C,\zeta})$ from below: Under $M_{j}^{\prime}=\frac{Cn\nu e^{-\nu}}{k}$, with respect to the randomness of $\mathbf{Z}$ we have $M_{j0}^{\prime}\sim\text{Bin}\big{(}\frac{Cne^{-\nu}\nu}{k},\rho\big{)}$, which should equal $\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}$ to ensure $j\in\mathcal{K}^{\prime}_{C,\zeta}$. Accordingly, for $1\leq j\leq k^{\xi}$, we proceed as follows:
$\mathbb{P}\big{(}j\in\mathcal{K}^{\prime}_{C,\zeta}\big{|}(\mathbf{T}_{s},\mathbf{T}_{1},...,\mathbf{T}_{j-1})\big{)}$ (E.12)
$\quad=\mathbb{P}\Big{(}M_{j}^{\prime}=\frac{Ce^{-\nu}\nu n}{k},M_{j0}^{\prime}=\frac{\zeta\cdot Cn\nu e^{-\nu}}{k}\big{|}(\mathbf{T}_{s},\mathbf{T}_{1},...,\mathbf{T}_{j-1})\Big{)}$ (E.13)
$\quad=\mathbb{P}\Big{(}M_{j}^{\prime}=\frac{Cn\nu e^{-\nu}}{k}\big{|}(\mathbf{T}_{s},\mathbf{T}_{1},...,\mathbf{T}_{j-1})\Big{)}\cdot\mathbb{P}\Big{(}\text{Bin}\Big{(}\frac{C\nu e^{-\nu}n}{k},\rho\Big{)}=\frac{\zeta\cdot C\nu e^{-\nu}n}{k}\Big{)}$ (E.14)
$\quad\geq\exp\left(-\frac{\nu ne^{-\nu}}{k}\Big{[}e^{\nu}D\big{(}Ce^{-\nu}\|e^{-\nu}\big{)}+C\cdot D(\zeta\|\rho)+o(1)\Big{]}\right)$ (E.15)
$\quad:=\exp\Big{(}-\frac{\nu ne^{-\nu}}{k}\big{[}f_{1}(C,\zeta,\rho,\nu)+o(1)\big{]}\Big{)}:=P_{\text{in}},$ (E.16)
where in (E.15) we substitute (E.11) and apply (A.4) in Lemma A.1 (with leading factor absorbed into the exponent), and in (E.16) we define
$f_{1}(C,\zeta,\rho,\nu)=e^{\nu}D\big{(}Ce^{-\nu}\|e^{-\nu}\big{)}+C\cdot D(\zeta\|\rho).$ (E.17)
@@
$k^{\prime}_{C,\zeta}=\sum_{j=1}^{k}\mathbbm{1}(j\in\mathcal{K}^{\prime}_{C,\zeta})\geq\sum_{j=1}^{k^{\xi}}\mathbbm{1}(j\in\mathcal{K}^{\prime}_{C,\zeta}):=k^{\prime}_{C,\zeta,\xi},$ (E.18)
and from (E.16) we have that $(k^{\prime}_{C,\zeta,\xi}|\mathbf{T}_{s})$ stochastically dominates $\text{Bin}(k^{\xi},P_{\text{in}})$, and thus
$\mathbb{P}\big{(}k^{\prime}_{C,\zeta,\xi}\leq z\,\big{|}\,\mathbf{T}_{s}\big{)}\leq\mathbb{P}\big{(}\text{Bin}(k^{\xi},P_{\text{in}})\leq z\big{)}$ (E.19)
for any $z$. Combining this property with (E.8) and (E.18), we can proceed as follows:
$\mathbb{P}\big{(}k_{C,\zeta}=0\,|\,\mathbf{T}_{s}\big{)}$ $\leq\mathbb{P}\Big{(}k_{C,\zeta}^{\prime}\leq C_{0}\log k\big{|}\mathbf{T}_{s}\Big{)}$ (E.20)
$\leq\mathbb{P}\Big{(}k_{C,\zeta,\xi}^{\prime}\leq C_{0}\log k\big{|}\mathbf{T}_{s}\Big{)}$ (E.21)
$\leq\mathbb{P}\Big{(}\text{Bin}(k^{\xi},P_{\text{in}})\leq C_{0}\log k\Big{)},$ (E.22)
@@
$\mathbb{P}\big{(}k_{C,\zeta}=0\big{)}\leq\mathbb{P}\Big{(}\text{Bin}(k^{\xi},P_{\text{in}})\leq C_{0}\log k\Big{)}.$ (E.23)
We now claim that to ensure that (C1) holds with $1-o(1)$ probability, it suffices to ensure that
$k^{\xi}P_{\text{in}}=\exp\Big{(}\xi\log k-\frac{\nu ne^{-\nu}[f_{1}(C,\zeta,\rho,\nu)+o(1)]}{k}\Big{)}\geq\exp\Big{(}\eta^{\prime}_{1}\log\frac{p}{k}\Big{)}\to\infty$ (E.24)
for some $\eta^{\prime}_{1}>0$. Indeed, if (E.24) holds, then by the Chernoff bound (see (A.2) in Lemma A.1), we have
$\mathbb{P}\Big{(}\text{Bin}(k^{\xi},P_{\text{in}})\leq C_{0}\log k\Big{)}$ $\leq\exp\Big{(}-k^{\xi}P_{\text{in}}\Big{[}\frac{C_{0}\log k}{k^{\xi}P_{\text{in}}}\log\Big{(}\frac{C_{0}\log k}{k^{\xi}P_{\text{in}}}\Big{)}+1-\frac{C_{0}\log k}{k^{\xi}P_{\text{in}}}\Big{]}\Big{)}$ (E.25)
$\leq\exp\Big{(}-\frac{1}{2}k^{\xi}P_{\text{in}}\Big{)},$ (E.26)
where the second line holds since $\frac{C_{0}\log k}{k^{\xi}P_{\text{in}}}\leq\frac{C_{0}\log k}{\exp(\eta^{\prime}_{1}\log\frac{p}{k})}=o(1)$ for large enough $k$. The claim thus follows by combining with (E.23).
It remains to deduce the threshold from (E.24). By $k=\Theta(p^{\theta})$ we have $\log k\sim\frac{\theta}{1-\theta}\log\frac{p}{k}$, and then by choosing $\xi$ arbitrarily close to $1$ we obtain (E.24) provided that the following holds for some $\eta_{1}>0$:
$n\leq(1-\eta_{1})\frac{\frac{\theta}{1-\theta}k\log\frac{p}{k}}{\nu e^{-\nu}f_{1}(C,\zeta,\rho,\nu)}.$ (E.27)
@@
The subsequent analysis is built on condition (C1), which ensures that $k_{C,\zeta}\geq 1$, and we consider a single specific index $j\in\mathcal{K}_{C,\zeta}$. We will study the placement of each non-defective item $j^{\prime}$ for $j^{\prime}\in\{k+1,\dotsc,p\}$. For the given $j\in\mathcal{K}_{C,\zeta}$ and a fixed $j^{\prime}\in\{k+1,\dotsc,p\}$, we define $R_{j^{\prime}}=|\mathcal{N}_{0,j^{\prime}}\cup\mathcal{M}_{j,j^{\prime}}|$ as the number of tests in $\mathcal{N}_{0}\cup\mathcal{M}_{j}$ that contain item $j^{\prime}$, and define $R_{j^{\prime}}^{\prime}$ as the number of placements of item $j^{\prime}$ that are in $\mathcal{N}_{0}\cup\mathcal{M}_{j}$. It is evident that $R_{j^{\prime}}^{\prime}\geq R_{j^{\prime}}$ always holds, and $R_{j^{\prime}}^{\prime}>R_{j^{\prime}}$ when item $j^{\prime}$ is placed more than once into some test in $\mathcal{N}_{0}\cup\mathcal{M}_{j}$. To formulate condition (C2), we are primarily interested in $R_{j^{\prime}}$, but it is more convenient to first study $R_{j^{\prime}}^{\prime}$. Recall that the event $\mathscr{A}_{3}$ in (E.5) states that
$ne^{-\nu}(1+o(1))=N_{0}\leq|\mathcal{N}_{0}\cup\mathcal{M}_{j}|\leq N_{0}+\ell\Delta=ne^{-\nu}(1+o(1)).$ (E.28)
Under this event, a placement of item $j^{\prime}$ increments $R_{j^{\prime}}^{\prime}$ by $1$ with probability $e^{-\nu}(1+o(1))$, and we thus have $R_{j^{\prime}}^{\prime}\sim\text{Bin}(\Delta,e^{-\nu}(1+o(1)))$ with respect to the randomness in placing item $j^{\prime}$ into tests.
Bounding the difference between $R_{j^{\prime}}$ and $R_{j^{\prime}}^{\prime}$: We are ultimately interested in $R_{j^{\prime}}$, but we have characterized $R_{j^{\prime}}^{\prime}$ as an intermediate step. Accordingly, we proceed to quantify the difference between $R_{j^{\prime}}$ and $R_{j^{\prime}}^{\prime}$. We envision that the $\Delta$ placements of item $j^{\prime}$ are done sequentially, and let $D_{j^{\prime}}$ be the number of “collisions”, defined as
$D_{j^{\prime}}=\sum_{i=1}^{\Delta}\mathbbm{1}\big{(}\text{the }i\text{-th placement coincides with the }i_{1}\text{-th placement for some }i_{1}<i\big{)}.$ (E.29)
@@
$R^{\prime}_{j^{\prime}}-D_{j^{\prime}}\leq R_{j^{\prime}}\leq R^{\prime}_{j^{\prime}},$ (E.30)
@@
$\mathbb{P}\Big{(}D_{j^{\prime}}\geq D_{0}\Big{)}\leq\binom{\Delta}{D_{0}}\Big{(}\frac{\Delta}{n}\Big{)}^{D_{0}}\leq\Big{(}\frac{e\Delta^{2}}{D_{0}n}\Big{)}^{D_{0}}.$ (E.31)
We take $D_{0}=\frac{10}{\theta}$ in (E.31) and apply a union bound over $j^{\prime}\in\{k+1,\dotsc,p\}$, yielding
$\mathbb{P}\Big{(}\max_{j^{\prime}\in\{k+1,\dotsc,p\}}D_{j^{\prime}}\geq\frac{10}{\theta}\Big{)}$ $\leq p\cdot\Big{(}\frac{e\theta\Delta^{2}}{10n}\Big{)}^{\frac{10}{\theta}}$ (E.32)
$=\exp\Big{(}\log p+\frac{10}{\theta}\log\Big{(}\frac{e\theta\nu^{2}n}{10k}\Big{)}-\frac{10}{\theta}\log k\Big{)}$ (E.33)
$\leq k^{-5},$ (E.34)
where (E.34) holds when $k$ is large enough because $p=\Theta(k^{1/\theta})$ and $n=\Theta(k\log k)$ (see Section A.2). In the subsequent analysis, we implicitly suppose (via a union bound) that we are on the event
$D_{j^{\prime}}<\frac{10}{\theta},\leavevmode\nobreak\ \forall j^{\prime}\in\{k+1,\dotsc,p\},$ (E.35)
@@
Studying the condition for (E.1): Given $R_{j^{\prime}}$, the randomness of the noise $\mathbf{Z}$ gives
$G_{j,j^{\prime},1}\sim\text{Bin}(R_{j^{\prime}},\rho),\leavevmode\nobreak\ \leavevmode\nobreak\ G_{j,j^{\prime},2}=R_{j^{\prime}}-G_{j,j^{\prime},1}.$ (E.36)
Thus, (E.1) can be expressed as $2G_{j,j^{\prime},1}-R_{j^{\prime}}>(1-2\zeta)\frac{Cn\nu e^{-\nu}}{k}$, or equivalently
$G_{j,j^{\prime},1}>\Big{(}\frac{1}{2}-\zeta\Big{)}\frac{Cn\nu e^{-\nu}}{k}+\frac{1}{2}R_{j^{\prime}}.$ (E.37)
@@
$\mathbb{P}\Big{(}G_{j,j^{\prime},1}>\big{(}\frac{1}{2}-\zeta\big{)}\frac{Cn\nu e^{-\nu}}{k}+\frac{1}{2}R_{j^{\prime}}\Big{)}$ (E.39)
$\geq\mathbb{P}\Big{(}R_{j^{\prime}}^{\prime}=\frac{dn\nu e^{-\nu}}{k}\Big{)}\mathbb{P}\Big{(}\text{Bin}(R_{j^{\prime}},\rho)>\big{(}\frac{1}{2}-\zeta\big{)}\frac{Cn\nu e^{-\nu}}{k}+\frac{1}{2}R_{j^{\prime}}\Big{|}R_{j^{\prime}}^{\prime}=\frac{dn\nu e^{-\nu}}{k}\Big{)}$ (E.40)
@@
$\quad\quad\quad\quad\cdot\mathbb{P}\Big{(}\text{Bin}\big{(}\frac{(1+o(1))dn\nu e^{-\nu}}{k},\rho\big{)}>\frac{d+C(1-2\zeta)}{2}\frac{n\nu e^{-\nu}}{k}\Big{)}$ (E.41)
$\geq\exp\left(-\frac{\nu ne^{-\nu}}{k}\Big{[}e^{\nu}\cdot D(de^{-\nu}\|e^{-\nu})+d\cdot D\Big{(}\frac{1}{2}+\frac{C(1-2\zeta)}{2d}\big{\|}\rho\Big{)}+o(1)\Big{]}\right)$ (E.42)
$:=\exp\Big{(}-\frac{\nu ne^{-\nu}}{k}\big{[}g(C,\zeta,d,\rho,\nu)+o(1)\big{]}\Big{)},$ (E.43)
where in (E.41) we use $R_{j^{\prime}}^{\prime}\sim\text{Bin}(\Delta,e^{-\nu}+o(1))$ along with the event (E.35) and $R_{j^{\prime}}^{\prime}=\frac{dn\nu e^{-\nu}}{k}$, from which (E.30) gives $R_{j^{\prime}}=\frac{dn\nu e^{-\nu}(1+o(1))}{k}$ (since $\frac{n}{k}\to\infty$ and $\frac{10}{\theta}$ is a constant); then in (E.42) we apply (A.4) in Lemma A.1 to the two probability terms (with leading factors absorbed into the exponent), and in (E.43) we introduce the following shorthand:
$g(C,\zeta,d,\rho,\nu)=e^{\nu}\cdot D\big{(}de^{-\nu}\|e^{-\nu}\big{)}+d\cdot D\Big{(}\frac{1}{2}+\frac{C(1-2\zeta)}{2d}\big{\|}\rho\Big{)}.$ (E.44)
@@
$g(C,\zeta,d,\rho,\nu)$ $=d\log d+(e^{\nu}-d)\log\frac{1-de^{-\nu}}{1-e^{-\nu}}+\frac{d+A}{2}\log\frac{d+A}{2d\rho}+\frac{d-A}{2}\log\frac{d-A}{2d(1-\rho)}.$ (E.45)
Optimizing $d$: Our next step is to specify $d$ that minimizes $g(C,\zeta,d,\rho,\nu)$ to render the tightest lower bound (E.43). Note that range of $d$ is given by $|A|\leq d\leq e^{\nu}$ in (E.38), and we define
$d^{*}=\text{arg}\min\leavevmode\nobreak\ g(C,\zeta,d,\rho,\nu),\leavevmode\nobreak\ \leavevmode\nobreak\ \text{subject to }|A|\leq d\leq e^{\nu}.$ (E.46)